OpenStudy (anonymous):

A steam catapult launches a jet aircraft (from rest) from the aircraft carrier John C. Stennis, giving it a speed of 200 mi/h in 3.00 s. (a) Find the average acceleration of the plane.

OpenStudy (stormfire1):

\[v_{final}=v _{initial}+at\]\[200mph = 0 + a * 3.00s\] Now just solve for a to get \[66.67 mph ^{2}\]

OpenStudy (anonymous):

v = u + at a = v-u/t a = 200 - 0/3 a = 66.7 mph^2